# appendixReferences

### v3

@@ 6.2 Social Networks and the Decision to Insure
We analyze an experiment that was originally described in Cai et al., (2015), quoting first a summary of the experiment from Choi, (2021):
@@
In an experiment described in Cai et al., (2015), rural farmers in China were randomly assigned to information sessions where they would be given the opportunity to purchase weather insurance. The sessions were randomized to give either high or low levels of information about the insurance product. First round sessions were held three days before second round sessions, so that first round attendees would have opportunity for informal conversations with their second round friends, in which they might share their opinions about the insurance product. Social network information was elicited, with the farmers instructed to list 5 close friends with whom they specifically discussed rice production or financial issues.
The goal of the experiment was to broadly demonstrate the importance of information sharing, by measuring its effects in a randomized setting. One of the conclusions of Cai et al., (2015) was that the decision to purchase insurance was affected not only by a farmer’s own treatment assignment, but also by that of their friends; specifically, farmers assigned to a second round low-information session were more likely to purchase insurance if more of their listed friends in the first round were assigned to a high-information session.
In the original analysis of the experiment, the unit outcomes were assumed to depend only on the information received by themselves and their listed friends. To justify this assumption, a subjective argument was made that the three day gap between information sessions was too short for information to travel more than one hop on the observed social network.
@@
In our analysis, we forgo such assumptions and estimate $\tau^{\textup{control}}$, the number of second round units whose outcome under direct control (i.e., being assigned to a low information session) would be affected by the assignment of first round farmers into high and low information sessions. To construct our estimators, we require knowledge of the randomization design, which is contained in Gao and Ding, (2023).
@@
Our point estimate equals 23%, which may be interpreted as meaning that at least 23% of second round farmers, if assigned to a low information session, would be affected by information given to the first round farmers (1-sided 95% CI: at least 9%). This point estimate of 23% is equal to the Hájek weighted comparison of second round units who received low information directly but had many first round friends with high information, versus those in the second round who received low information directly and had few or no first round friends with high information:
$\hat{\tau}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1\{X_{i}=0,W_{i}=1\}N}{P(X_{i}=0,W_{i}=1)\hat{N}_{1}}Y_{i}-\frac{1\{X_{i}=0,W_{i}=0\}N}{P(X_{i}=0,W_{i}=0)\hat{N}_{0}}Y_{i}\right)$
@@
where $i\in[N]$ enumerates the second round units, $X_{i}=0$ if unit $i$ was assigned to a low information session, $W_{i}=1$ if all of unit $i$’s first round friends received high information, $Y_{i}$ denotes unit $i$’s decision of whether or not to purchase insurance, and $\hat{N}_{a}=\sum_{i=1}^{N}1\{X_{i}=0,W_{i}=a\}/P(X_{i}=0,W_{i}=a)$ for $a\in\{0,1\}$.
@@
For comparison, we consider an EATE-type treatment effect that considers the relative effects of receiving $(X_{i},W_{i})$ equal to $(0,1)$ versus $(0,0)$:
treatment $\textup{treatment effect}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=1]-\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=0]\right),$
where the expectation is taken over the randomization of treatment under the experiment design. For this target parameter, the estimation method of Gao and Ding, (2023) gives a Hájek-normalized point estimate of 23%, and 95% CI of $[2\%,45\%]$. This confidence interval requires an assumption of no interference between villages. Such an assumption might be debatable, as farmers listed cross-village friendships. In contrast, no assumptions on interference are required for our estimand. Thus for the purposes of demonstrating the presence of social influence, our estimand may be an appropriate target parameter, and has tighter, less questionable CIs when compared to an analogous treatment effect.
Appendix
Appendix A Proofs
The organization of this section is as follows. In Section A.1, we present helper lemmas that will be used in our proofs. Section A.2 proves Theorems 1 - 6. Finally, in Section A.3 we prove Lemmas 1 - 9.
A.1 Auxilliary Results

### v4

@@ 6.2 Social Networks and the Decision to Insure
We analyze an experiment that was originally described in Cai et al., (2015), and re-analyzed in Gao and Ding, (2023).666data available at https://www.openicpsr.org/openicpsr/project/113593/version/V1/view In this experiment, rural farmers were randomized to receive high or low information regarding a crop insurance product, with opportunity to purchase the product at the end of the information session. Before attending their assigned sessions, the farmers could communicate informally with others who had attended earlier sessions (sessions were administered in two rounds, with a three day gap between them), allowing for sharing of information through social contacts. Social network information was elicited, with the farmers instructed to list 5 close friends with whom they specifically discussed rice production or financial issues. Farmers assigned to a second round low-information session were more likely to purchase insurance if more of their listed friends in the first round were assigned to a high-information session.
Table 8 in the supplement shows results from the original regression analysis of Cai et al., (2015), which found that for farmers assigned to to low information sessions, each friend assigned to a first round high-information session caused an 8.9% increase in insurance purchasing rates (95% CI: [4.5%, 13.2%]). In their analysis, it was assumed that a unit’s outcome depends only on the information received by themselves and their listed friends. To justify this assumption, a subjective argument was made that the three day gap between information sessions was long enough for information to reach direct friends, but too short for information to diffuse across an entire village.
@@
In our analysis, we make no assumptions on interference and estimate $\tau^{\textup{control}}$, the number of second round units whose outcome under direct control (i.e., assignment to a low info session) would be affected by the assignment of first round farmers into high and low information sessions. Interference is permitted between distant or even disconnected units; for example, a farmer might ask a question during an information session, affecting the information received by (and subsequently shared by) all farmers at the session. To construct our estimators, we require knowledge of the randomization design, which used stratification and is described in Gao and Ding, (2023).
@@
Our asymptotic lower bound equals 23%, which may be interpreted as meaning that at least 23% of the 880 second round farmers, if assigned to a low information session, would be affected by information given to the first round farmers (1-sided 95% CI: at least 9%). This point estimate of 23% is equal to a Hájek weighted comparison of the second round units receiving low information, grouped according to their number of friends receiving high information in the previous round:
$\frac{\Delta^{\textup{Haj}}}{N}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1\{X_{i}=0,W_{i}=1\}N}{P(X_{i}=0,W_{i}=1)\hat{N}_{1}}-\frac{1\{X_{i}=0,W_{i}=0\}N}{P(X_{i}=0,W_{i}=0)\hat{N}_{0}}\right)Y_{i}$ (48)
@@
Here $i\in[N]$ enumerates the $N=880$ second round units with at least one first round friend; $X_{i}=0$ indicates that unit $i$ was assigned to a low information session; $W_{i}=1$ indicates that all of unit $i$’s first round friends received high information; $Y_{i}=1$ indicates that unit $i$ purchased insurance; and $\hat{N}_{a}=\sum_{i=1}^{N}1\{X_{i}=0,W_{i}=a\}/P(X_{i}=0,W_{i}=a)$ for $a\in\{0,1\}$. Additional results can be found in Appendix B of the supplemental materials.
Table 7 in the supplement shows results from Gao and Ding, (2023), who estimate the spillover effects of having friends receive high information. They report weakly significant or non-significant spillover effects of the first round assignments on the second round units, assuming conditions on interference which are technical in nature. It may be difficult to judge whether these conditions are suitable for this application. In contrast, no assumptions on interference are required for our estimates. For the purposes of demonstrating the presence of social influence, our estimand may be an appropriate target parameter, and has tighter, less questionable CIs. Thus it may serve as a useful complement to prior studies of this experiment.
Supplemental Materials
The supplement contains discussion of future work; further discussion and results for the simulations and data example; and proofs of all propositions and theorems, with generalization of Theorems 1 - 3 to include two-stage randomization.
Supplementary Material for “Agnostic Characterization of Interference in Randomized Experiments”
Appendix A Directions for Future Work
In principle, our approach can be extended to settings with categorical or continuous outcomes in various ways. The simplest approach is simply to binarize the outcomes, requiring a choice of threshold. Under binarization, our estimates will continue to lower bound the quantity of interest. For example, $\tau^{\textup{basic}}$ after outcome binarization denotes the number of units whose binarized outcome is affected by treatment, which is a subset of the units whose original outcome (i.e., pre-binarization) is affected by treatment.
To avoid thresholding of continuous-valued outcomes, we conjecture that rank-based estimators reminiscent of Rosenbaum, (2007) may also be devised, or kernel-based density estimates of the outcomes conditional on receiving treatment or control. Under the latter approach, the total variation difference between the conditional densities takes the place of the diff-in-means contrast $\Delta$, although we are unsure whether the optimization approach of (46) can be successfully adapted for inference.
We expect that our approach can be adapted to most experiment designs that have been considered for estimation of treatment effects under the assumption of a correctly specified exposure model. This is because consistency and normality theorems that require a correctly specified exposure model (normally a strong assumption) can be applied to $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ without question, as these estimators only consider units in $\mathcal{I}$}, for which the exposure model condition holds by definition.
Experiments involving group randomization or random edge formation may be of particular interest. As previously mentioned, for such experiments $\tau^{\textup{basic}}$ characterizes interference, as it denotes the number of units who are affected by their randomly assigned neighbors or group members. A possible approach might be similar to Li et al., (2019) which studied peer effects, while removing assumptions regarding interference that were required by that work. Let $A\in\{0,1\}^{N\times N}$ denote a random adjacency matrix whose distribution is jointly exchangeable following Bickel and Chen, (2009), so that $A$ and $\Pi A\Pi^{T}$ are equally probable for any permutation matrix $\Pi\in\{0,1\}^{N\times N}$. For group randomization experiments, the adjacency matrix $A$ would represent the randomly assigned groups as a collection of cliques, while edge formation experiments could use various network models, such as latent space or stochastic blockmodel (Bickel and Chen, , 2009; Hoff et al., , 2002). Let $O=(O_{1},\ldots,O_{N})\in\{0,1\}^{N}$ denote binary pre-treatment covariates for the $N$ units, with covariate classes given by $\mathcal{O}_{a}=\{i:O_{i}=a\}$ for $a\in\{0,1\}$. Let the treatment variables $X=(X_{1},\ldots,X_{N})$ denote the thresholded number of neighbors whose observable level is 1, so that
$X_{i}=1\left\{\sum_{j=1}^{N}A_{ij}O_{j}\geq t_{O_{i}}\right\}$
where $t_{0}$ and $t_{1}$ denote thresholds for units in $\mathcal{O}_{0}$ and $\mathcal{O}_{1}$ respectively. We conjecture that conditional on the number of units in $\mathcal{O}_{0}$ and $\mathcal{O}_{1}$ who were treated (i.e., for whom $X_{i}=1$), inference will take simple form in which the conditional distribution of $X$ is given by sampling without replacement within $\mathcal{O}_{k}$ for $k=0,1$. We additionally conjecture that this property will facilitate adaptation of the necessary asymptotic results, given by Theorems 1 - 3, to this setting.
Appendix B Additional Notes on Examples
Simulations
Choi, (2023) gives a proof-of-concept attempt to estimate $\tau^{\textup{indirect}}$, which was tested on the simulations described in Section 6.1. The approach was found to be lacking in power compared to ours, with 1-sided intervals whose boundaries were considerably closer to zero (and included zero considerably more often), as summarized in Table 4.
Data Example
For the estimation of $\tau^{\textup{control}}$ in the data analysis, $W_{i}$ was defined to equal 1 if all of unit $i$’s first round friends were assigned to high info sessions. This choice of threshold for $W_{i}$ was inspired by Choi, (2024), who found an effect threshold at 2 treated (i.e., first round high info) friends. Thresholding $W_{i}$ at 2 treated friends gives similar results for $\tau^{\textup{control}}$ (point estimate: 22%, 1-sided 95% CI: at least 8.8%), but restricts analysis to units with at least 2 first round friends. Thresholds exceeding 2 would remove most of the units from the analysis. These results are summarized in Table 5, which shows estimates of $\tau^{\textup{control}}$ for varying choices of threshold and under Bonferoni correction (which did not cause the intervals to become weakly significant or non-significant).
Choi, (2023) Our approach
NN P(Significant) Avg. $N^{-1}LB_{1-\alpha}$ bias P(Significant) Avg. $N^{-1}LB_{1-\alpha}$ bias
$N=250$ 5% 0.0005 -0.26 80% 0.076 -0.15
$N=500$ 20% 0.0027 -0.23 97% 0.093 -0.13
Table 4: Comparison of Choi, (2023) and our approach for estimating $\tau^{\textup{indirect}}$ on simulated data. P(Significant) denotes the fraction of trials for which the 1-sided interval excluded zero. Avg. $N^{-1}LB_{1-\alpha}$ denotes the average value of the lower boundary (as a fracton of $N$) of the 1-sided interval. Bias reports the average difference (as a fraction of $N$) between the point estimator and $\tau^{\textup{indirect}}$.
Threshold $t_{i}$ Point Estimate 1-sided 95% CI Bonferoni corrected
$\geq 1$ treated friend 8.8% at least 0% at least 0%
$\geq 2$ treated friends 22% at least 8.8% at least 5.7%
All 1st round friends treated 23% at least 9.9% at least 7.0%
Table 5: Estimates of $\tau^{\textup{control}}$ for data example of Cai et al., (2015), under varying choices of threshold $t_{i}$. Bonferoni correction is for 3 tests ($\approx 98.33\%$ confidence level).
Our result is significantly stronger than the proof-of-concept attempt of Choi, (2024) to estimate the number of second round units who were affected by the first round treatments. Their analysis did not produce a point estimate, and gave a much more conservative value of 6 units (or 2% of a matched subset of units) for the boundary of the 1-sided confidence interval, which is smaller than ours by a factor of 13. The approach of Choi, (2023) also performs poorly, producing a conservative point estimate of at least 6 affected units, which was not statistically significant (95% CI: at least zero units).
@@ Appendix B Additional Notes on Examples
We may also compare to an EATE-type treatment effect that considers the relative effects of receiving $(X_{i},W_{i})$ equal to $(0,1)$ versus $(0,0)$:
$\textup{treatment effect}=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=1]-\mathbb{E}[Y_{i}|X_{i}=0,W_{i}=0]\right),$ (49)
where the expectation is taken over the randomization of treatment under the experiment design. For this target parameter, the estimation method of Gao and Ding, (2023) gives a Hájek-normalized point estimate of 24%, with a 95% CI ranging in width from $1.4\%,43\%]$ to $5.3\%,40\%]$ depending on the choice of kernel matrix $K_{n}$, as summarized in Table 6.
The role of the kernel matrix $K_{n}$ is to choose which pairs of units should be used by the variance estimator of Eq. (11) in Gao and Ding, (2023). Given an observed network, they propose a bandwidth rule to balance between bias (from excluding important pairs of units whose residuals are strongly correlated) and variance (from including too many unit pairs). In this rule, pairs of units are used for variance estimation if the length of the shortest path between them is less than or equal to a parameter $b_{n}$. However, it seems difficult to discern when one choice of $b_{n}$ will be more accurate than another, or if any of the tested choices will necessarily be accurate for the fixed size of the data at hand. As a result, the correctness of our intervals for (49) may be questionable.
Table 7 shows selected results from the analysis of Gao and Ding, (2023). It shows point estimates and estimated standard errors for the effect on insurance takeup of having at least one first round friend assigned to a high information session. Under all choices of $b_{n}$ exceeding zero, the effect is not significant at 95% confidence level. [^7] Gao and Ding, (2023) also report weak significance using uncorrected variance estimates (which may be anticonservative), in Tables S4 and S5 of their paper. Unlike our analysis, Gao and Ding, (2023) threshold the number of treated friends at 1, and includes second round units who were assigned to high info sessions. Changing both of these choices yields the above-mentioned effect of (49).
Table 8 shows selected regression results from the original analysis of Cai et al., (2015), which we include for reference.
Software
We used the commerical solver Gurobi, through MATLAB after installing the optimization toolbox YALMIP (Lofberg, , 2004).
Choice of $K_{n}$ (i.e., criteria for $K_{n}(i,j)=1$ to hold) 95% CI
$b_{n}=2$ (Units $i$ and $j$ have path distance $\leq 2$): $1.7\%\,,43\%]$
$b_{n}=4$ (Units $i$ and $j$ have path distance $\leq 4$): $2.8\%\,,42\%]$
Units $i$ and $j$ in the same village, or have friends
      in the same village: $5.3\%\,,40\%]$
Units $i$ and $j$ in same strata, or have friends
      in the same strata: $1.4\%\,,43\%]$
Table 6: 95% CI for EATE estimate of (49), using Eq. (10) of Gao and Ding, (2023) under varying choices of kernel matrix $K_{n}$. The last choice of $K_{n}$ (same strata/friends in same strata) captures dependencies that arise when treatments are assigned by stratified random sampling without replacement, or equivalently when inference is conditional on the stratified treatment rates.
Unadj Add Sat
Point Estimate: 0.056 0.058 0.055
std. err, $b_{n}=0$: 0.027 0.027 0.027
std. err, $b_{n}=2$: 0.033 0.033 0.033
std. err, $b_{n}=3$: 0.031 0.031 0.031
std. err, $b_{n}=4$: 0.031 0.031 0.031
std. err, $b_{n}=5$: 0.032 0.032 0.032
Table 7: Estimated spillover effects reported by Table S5 of Gao and Ding, (2023), for “two-dimensional exposure mapping” model and “WLS+ SE” standard errors. Regression specifications, where $z_{i}$ and $x_{i}$ respectively denote treatment and covariate variables, are as follows. Unaj: $Y_{i}\sim 1+z_{i}$. Add: $Y_{i}\sim 1+z_{i}+x_{i}$. Sat: $1+z_{i}+(x_{i}-\bar{x})+z_{i}(x_{i}-\bar{x})$.
Regression Variable (2) (4)
Invited to high info session: 0.0298 0.0809
(0.0332) (0.0397)
Fraction of friends invited to 1st round high info session (NET): 0.291*** 0.444***
(0.0820) (0.109)
(NET)*(invited to high info session): -0.329**
Table 8: Regression results reported by Table 2 of Cai et al., (2015) for specifications (2) and (4). For the effect of treating a single friend, Cai et al., (2015) divided the NET coefficient by 5, the number of elicited friends.
Appendix C Asymptotic Results
The organization of this section is as follows. In Section C.1, we present Assumption 2, a more general version of Assumption 1 which allows other experiment designs to be considered. We also present Theorems 4 and 5 which respectively describe conditions under which Bernoulli randomization designs and two-stage randomization designs will satisfy the assumption. In Section C.2, we present Theorems 68, which are generalized versions of Theorems 13, and respectively give results on consistency, variance estimation, and interval coverage under Assumption 2. Proofs of all propositions and theorems are given in Section D.
C.1 Experiment Designs
Assumption 2 describes a class of experiment designs, which includes independent Bernoulli randomization and two stage randomization designs. It assumes that units can be divided into $M$ groups $G_{1},\ldots,G_{M}$, such that the subvectors $X_{G_{1}},\ldots,X_{G_{M}}$ of $X$ are independent random variables.
Assumption 2.
The units may be partitioned into $M$ non-overlapping groups $G_{1},\ldots,G_{M}\subset[N]$ such that the subvectors $X_{G_{1}},\ldots,X_{G_{M}}$ are independent random variables, and the following hold:
  1. 1.
    For fixed constants $C>0$ and $0<\delta\leq 1$ it holds as $N\rightarrow\infty$ that
    $|G_{m}|\leq C\frac{N}{M}\qquad\textup{ and }\qquad|G_{m}|\leq N^{1-\delta}\qquad\forall m\in[M],$ (50)
  2. 2.
    There exists group neighborhoods $\eta_{1},\ldots,\eta_{M}\subset[M]$ satisfying for fixed $d>0$ that
    $\max_{m}|\eta_{m}|$ $\leq d$ (51)
    $\max_{\ell}\sum_{m=1}^{M}1\{\ell\in\eta_{m}\}$ $\leq d$ (52)
    such that the design neighborhoods $\{\Gamma_{i}\}_{i=1}^{N}$ satisfy
    $\Gamma_{i}=\bigcup_{m\in\eta_{g(i)}}G_{m},$ (53)
    where $g(i)$ denotes the group of unit $i$ (that is, $g(i)=m$ if $i\in G_{m}$).
  3. 3.
    For a fixed constant $\rho>0$ it holds for all $i,j\in[N]$ and $a,b\in\{0,1\}$ that
    $P(\bar{Z}_{i}=a|T_{i}=t_{i})$ $\geq\rho$ $\forall t_{i}\in\mathcal{T}_{i}$ (54)
    $\frac{P(T_{i}=t_{i})P(T_{j}=t_{j})}{P(T_{i}=t_{i},T_{j}=t_{j})}$ $\leq\frac{1}{\rho}$ $\forall(t_{i},t_{j})\in\mathcal{T}_{ij}$ (55)
    $P(\bar{Z_{i}}=a,\bar{Z}_{j}=b|T_{i}=t_{i},T_{j}=t_{j})$ $\in\{0\}\cup[\rho,1]$ $\forall(t_{i},t_{j})\in\mathcal{T}_{ij}$ (56)
    where $\mathcal{T}_{i}$ denotes the support of $t_{i}$, and $\mathcal{T}_{ij}$ denotes the support of $(t_{i},t_{j})$
The conditions of Assumption 2 may be interpreted as follows. Condition (50) requires the size of the largest group to be within a constant factor of the average group size, and allows the group sizes to grow sublinearly as $N\rightarrow\infty$. Condition (53) restricts the design neighbornoods $\{\Gamma_{i}\}_{i=1}^{N}$ to equal equal each unit’s own and neighboring groups, where the group neighborhoods $\{\eta_{m}\}_{m=1}^{M}$ are required by (51)-(52) to have bounded size and overlap. Conditions (54)-(56) are support conditions, which bound the magnitude of the inverse probability weights used by $\hat{\tau}_{k}^{\textup{Haj}}$ and $\hat{V}_{k}$.
Example: Bernoulli Design
Theorem 4 states that the setting of Assumption 1, in which the treatments $\{X_{i}\}_{i=1}^{N}$ are independent Bernoulli random variables, can be modeled under Assumption 2 by assigning each unit to its own group, so that $M=N$.
Theorem 4.
Let Assumption 1 hold. Then Assumption 2 holds with $M=N$.
Example: Two-stage Design
Theorem 5 gives conditions under which $|\Gamma_{i}|$ may grow at rate $N^{1-\delta}$ for any $\delta>0$, in which the treatments are randomized by a two-stage design. In this setting, the units are divided into groups, and treatment is randomized first at the group level, and then at the unit level conditional on group level treatment. The contrast variable $\bar{Z}_{i}$ may consider each unit’s direct treatment, the treatment of their close associates $\mathcal{N}_{i}$, and the number of treated units in their own and neighboring groups, where the group sizes may grow with $N$ at any sublinear polynomial rate.
Theorem 5.
Let $g(i)$ denote the group membership of unit $i$, so that $g(i)=m$ if $i\in G_{m}$. Assume the following hold for fixed $C>0$, $0<\delta\leq 1$, $d>0$, and $p>0$ as $N\rightarrow\infty$
  1. A1.
    The groups $G_{1},\ldots,G_{M}$ form a partition of $[N]$ satisfying (50), with group neighborhoods $\{\eta_{m}\}_{m=1}^{M}$ which satisfy the size and overlap conditions (51) and (52).
  2. A2.
    Each treatment subvector $X_{G_{m}}$ is independently generated by sampling $\theta_{m}$ units without replacement from $G_{m}$, where $\theta_{1},\ldots,\theta_{M}$ are independent random variables with support satisfying
    $\operatorname{supp}(\theta_{m})\subseteq\{0\}\cup\{\theta_{m}^{\textup{low}},\ldots,\theta_{m}^{\textup{high}}\}\cup\{|G_{m}|\}$ (57)
    for $\theta_{m}^{\textup{low}}\geq p|G_{m}|$ and $\theta_{m}^{\textup{high}}\leq(1-p)|G_{m}|$, and which additionally satisfy
    $P(\theta_{m}=t)\geq p\qquad\forall t\in\operatorname{supp}(\theta_{m})$ (58)
  3. A3.
    For $i\in[N]$, $\bar{Z}_{i}$ satisfies
    $\bar{Z}_{i}$ $=\zeta_{i}(X_{\mathcal{N}_{i}},\theta_{\eta_{g(i)}}),$ $\forall i\in[N],$ (59)
    for some set of mappings $\{\zeta_{i}\}_{i=1}^{N}$ and unit neighborhoods $\{\mathcal{N}_{i}\}_{i=1}^{N}$. Each neighborhood $\mathcal{N}_{i}$ is a subset of unit $i$’s own and neighboring groups so that
    $\mathcal{N}_{i}\subset\bigcup_{m\in\eta_{g(i)}}G_{m}$ (60)
    holds, and also satisfies $i\in\mathcal{N}_{i}$ and the size and overlap conditions
    $\max_{i}|\mathcal{N}_{i}|\leq d\qquad\textup{ and }\qquad\max_{i}\sum_{j=1}^{N}1\{i\in\mathcal{N}_{j}\}\leq d$ (61)
  4. A4.
    Either $T_{i}=1$ for all $i\in[N]$, or $T_{i}=X_{i}$ for all $i\in[N]$, or $T_{i}=X_{\mathcal{N}_{i}}$ for all $i\in[N]$
  5. A5.